OpenStudy (anonymous):

solve 4x+5y=9 2x+3y=3

OpenStudy (jiteshmeghwal9):

First equal one of the variables in both equations 1(4x+5y)=1(9) 2(2x+3y)=2(3) --------------- 4x+5y=9 -4x-6y=-6 ----------- -y=3 y=-3 Now put this value in one of the equation to gt x :)

OpenStudy (anonymous):

thank you

OpenStudy (jiteshmeghwal9):

yw :)

OpenStudy (anonymous):

-6y/6 = 1 @jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

multiply 6 both sides & then divide -6 both sides

OpenStudy (anonymous):

ok